ERRATUM: PERIODIC W -GRAPHS

G. LUSZTIG

I thank Jens C. Jantzen for pointing out the following misprints. In the second line of the display in 3.2, replace s ∉ L by s ∈ L. In 3.8, replace the last H by M in the line: ”It is clear that M(U) is an H-

submodule of H”. In 14.2, in the third line of the first display, replace vA_{a,b} by vA_{a}. In the line 14.3(c) replace v^{−1}Z^{−1} by v^{−1}Z[v^{−1}].